OpenStudy (anonymous):

write the equation of the line in point slope form. has a slope of 5/2 and passes through point (1,8) ?

OpenStudy (anonymous):

@jim_thompson5910 ??

OpenStudy (anonymous):

idk what the answer is lol @jim_thompson5910

OpenStudy (anonymous):

idk how to do that @jim_thompson5910

OpenStudy (anonymous):

5/2 @jim_thompson5910

OpenStudy (anonymous):

idk divide by 8?? @jim_thompson5910

OpenStudy (anonymous):

5/2 +b =8?? @jim_thompson5910

OpenStudy (anonymous):

dived by 3 and can we speed this up @jim_thompson5910

OpenStudy (anonymous):

subtraction

OpenStudy (anonymous):

subtract 5 from both sides

OpenStudy (anonymous):

3

OpenStudy (anonymous):

subtract 5/2 from boyj sides

OpenStudy (anonymous):

-5.5

OpenStudy (anonymous):

1

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

idk what is it? @jim_thompson5910

OpenStudy (anonymous):

@Ahmad1 ??

OpenStudy (anonymous):

idk what its equal to

OpenStudy (anonymous):

6/2??

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

mmmmmmmmmm what part you did not get so far @broderick365

OpenStudy (anonymous):

everything @Ahmad1

OpenStudy (anonymous):

@jim_thompson5910 is no help

OpenStudy (anonymous):

your welcome!

OpenStudy (anonymous):

if you have y=mx+b and m is the slope which is 5/2 , then the equation becomes y=5x/2+b just substitute the point you know which is (1,8) instead of (x,y) then 8=(5/2)*1+b b=8-5/2 then b= 11/2,,,, so we have b and now .. the equation is y=(5/2)x+(11/2)

OpenStudy (anonymous):

y - y1 = m(x - x1) slope(m) = 5/2 (1,8) x1 = 1 and y1 = 8 now just sub that into the formula

OpenStudy (anonymous):

y - 8 = 5/2(x - 1) y - 8 = 5/2x - 5/2 y = 5/2x - 5/2 + 8 y = 5/2x - 5/2 + 16/2 y = 5/2x + 11/2

OpenStudy (anonymous):

is that in point slope form?? @kelliegirl33 and @Ahmad1 ??

OpenStudy (anonymous):

oh....point slope form......the answer is : y - 8 = 5/2(x - 1)

OpenStudy (anonymous):

that's correct, just use the formula y-y1 = m(x-x1)